# Leveraging chaos in the training of artificial neural networks

### v1

@@ III Onset of sensitivity to initial conditions and training efficiency: Supplementary results
Figure S6: Lyapunov exponent (left column), percentage of valid exponents ($\rho$) and mean convergence time ($\langle\tau\rangle$) as functions of the learning rate ($\eta$). Results for the MNIST task with a single hidden layer and tanh activation function, now including $L_{2}$-regularization.
Finally we present the results obtained for the CIFAR-10 task in Figure S7. Although the prediction task is not successfully solved, since it is not the main focus of our study, we include this case to demonstrate that the observed phenomenon persists. Networks with this type of topology (MLP) typically do not achieve accuracies higher than 50% on CIFAR-10. Achieving accuracies of approximately 70% usually requires more sophisticated MLP architectures, such as including the use of linear bottleneck layers [45]. For this reason, we lower the accuracy threshold to 30% and observe that the same characteristic behaviour is maintained.
@@
Figure S7: Lyapunov exponent (left column), percentage of valid exponents ($\rho$) and mean convergence time ($\langle\tau\rangle$) as functions of the learning rate ($\eta$) for each activation function. Results for the CIFAR-10 task with a target accuracy of 30%.
Evolution of sharpness
@@
Following the approach in [38], we measure the evolution of sharpness during training, defined as the maximum eigenvalue of the Hessian of the training loss with respect to the model parameters. This measure characterises the local curvature of the loss surface and provides insight into the dynamics of the optimisation process. The sharpness is tracked across epochs for the model trained with the MNIST dataset and tanh as activation function.
@@
We use power iteration to approximate the top eigenvalue of the Hessian. This method is computationally efficient and does not require explicit computation and storage of the full Hessian matrix, so it is feasible to apply it during training. We track the evolution of the sharpness over training by measuring the top eigenvalue of the Hessian matrix of the loss function, as shown in Figure S8. For learning rates that minimize $\langle\tau\rangle$, we find that, in line with theoretical considerations [38], sharpness remains bounded once the training has stabilised and closely matches the theoretical limit of $2/\eta$, as indicated by the red dashed line. The results shown correspond to learning rates $\eta=7.5$ and $\eta=10$.
@@
Figure S8: Evolution of sharpness as a function of training epochs. The theoretical value $2/\eta$ is indicated by the dashed red line for reference. Results are shown for learning rates values $\eta=7.5$ and $\eta=10$.

### v2

@@ Abstract
Figure 11: Lyapunov exponent (left column), percentage of valid exponents ($\rho$) and mean convergence time ($\langle\tau\rangle$) as functions of the learning rate ($\eta$). Results for the MNIST task with a single hidden layer and tanh activation function, now including $L_{2}$-regularization.
Finally, we consider a convolutional architecture for MNIST consisting of two convolutional layers (8 and 16 kernels, $3\times 3$, padding=1), each follow by $2\times 2$ max pooling, and a fully connected output layer (9,098 trainable parameters). Convolutional layers use ReLU activations and weights are initialized from $\mathcal{N}(0,0.01)$. Training is performed for up to 2,000 epochs with 40 initial conditions and a target accuracy of 80%. As shown in Fig. 12, the characteristic sweet spot in learning rate remains visible.
@@
Figure 12: Lyapunov exponent (left column), percentage of valid exponents ($\rho$) and mean convergence time ($\langle\tau\rangle$) as functions of the learning rate ($\eta$). Results correspond to the CNN applied to the MNIST classification task.
We conclude with the CIFAR-10 dataset [52] (50,000/10,000 train/test samples; $\times32$ RGB images) using the 3072–256–10 MLP architecture (789,258 parameters). Since MLPs typically achieve limited performance on this task, we set the target accuracy to 30%. As shown in Fig. 13, the same qualitative behavior is observed despite the increased task complexity.
Figure 13: Lyapunov exponent (left column), percentage of valid exponents ($\rho$) and mean convergence time ($\langle\tau\rangle$) as functions of the learning rate ($\eta$) for each activation function. Results for the CIFAR-10 task with a target accuracy of 30%.
@@
Appendix C: Evolution of sharpness. Following the approach in [41], we measure the evolution of sharpness during training, defined as the maximum eigenvalue of the Hessian of the training loss with respect to the model parameters. This measure characterises the local curvature of the loss surface and provides insight into the dynamics of the optimisation process. The sharpness is tracked across epochs for the model trained with the MNIST dataset and tanh as activation function.
@@
We use power iteration to approximate the top eigenvalue of the Hessian. This method is computationally efficient and does not require explicit computation and storage of the full Hessian matrix, so it is feasible to apply it during training. We track the evolution of the sharpness over training by measuring the top eigenvalue of the Hessian matrix of the loss function, as shown in Figure 14. For learning rates that minimize $\langle\tau\rangle$, we find that, in line with theoretical considerations [41], sharpness remains bounded once the training has stabilised and closely matches the theoretical limit of $2/\eta$, as indicated by the red dashed line. The results shown correspond to learning rates $\eta=7.5$ and $\eta=10$.
@@
Figure 14: Evolution of sharpness as a function of training epochs. The theoretical value $2/\eta$ is indicated by the dashed red line for reference. Results are shown for learning rates values $\eta=7.5$ and $\eta=10$.